In an analysis of variance problem involving 3 treatments and 10 observations per treatment, SSE = 399.6. The MSE for this situation is which?

133.2

13.32

14.8

30.0
Answer #1

Given that, number of treatments = k = 3

number of observations = n = 10

Degrees of freedom for error = k(n - 1) = 3 * (10 - 1) = 27

SSE = 399.6

We want to find, MSE

MSE = 399.6 / 27 = 14.8

Answer : 14.8